A car move at a constant velocity of 20m/s for 10seconds and then accelerated uniformly to a velocity of 30m/s over 15 seconds. Calculate the distance covered during the motion
1
Expert's answer
2021-06-14T16:17:09-0400
"s_1=v_1t_1"

"s_1=20\ m/s\cdot 10s=200\ m"

"v_2=v_1+at_2"

"a=\dfrac{v_2-v_1}{t_2}"

"s_2=v_1t_2+\dfrac{at_2^2}{2}=v_1 t_2+\dfrac{(v_2-v_1)t_2}{2}"

"s_2=20 \ m/s\cdot15\ s+\dfrac{(30 \ m/s-20 \ m/s)15\ s}{2}=375\ m"


"s=s_1+s_2"

"s=200\ m+375\ m=575\ m"

The distance covered during the motion is 575 metres.